OpenStudy (anonymous):

What is the solution of log3(x + 7) + log33 = 2? (Points : 1) 4 5 10 no solution

Nnesha (nnesha):

addition change to ??

OpenStudy (elonasushchik):

3(x+7)=3^2=9 => x+7 = 3 => x= - 4 so negative 4 is the answer. Medal and fan thx!!

Nnesha (nnesha):

@ElonaSushchik http://openstudy.com/code-of-conduct welcome to open study!!

OpenStudy (elonasushchik):

@brittney_fraley did you understand how I explained it to you?..